OpenStudy (anonymous):

find all the solutions to the equation 2sin theta-1=0 write it in terms of radians

OpenStudy (anonymous):

2 sin(theta)-1=0 or 2 sin(theta-1)=0

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

I will assume 2 sin(theta)=1 sin(theta)=1/2 Now let's think, Sin of what angle gives you 1/2

OpenStudy (anonymous):

Remember unit circle

OpenStudy (anonymous):

pi/6

OpenStudy (anonymous):

yes, what else?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

sin(150)=sin(5pi/6)=1/2

OpenStudy (anonymous):

But that's not all because if you go around in full circle pi/6+2Pi (30+360=390) you are right where you were so same value of Sin() So it is not just pi/6. pi/6, pi/6+2pi, pi/6+2pi+2pi+..... pi/6,pi/6+2pi, pi/6+4pi +........ = pi/6+2n*pi

OpenStudy (anonymous):

Do the same with sin(150)=sin(5pi/6)=1/2